Titan Mining Corporation has 7.2 million shares of common stock outstanding, 260,000 shares of 4.4 percent preferred stock outstanding, and 145,000 bonds with a semiannual coupon rate of 5.5 percent outstanding, par value $2,000 each. The common stock currently sells for $65 per share and has a beta of 1.15, the preferred stock has a par value of $100 and currently sells for $91 per share, and the bonds have 15 years to maturity and sell for 105 percent of par. The market risk premium is 7.5 percent, T-bills are yielding 3.1 percent, and the company’s tax rate is 21 percent.

a.

What is the firm’s market value capital structure? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., .1616.)

b. If the company is evaluating a new investment project that has the same risk as the firm’s typical project, what rate should the firm use to discount the project’s cash flows? (Do not round intermediate calculations enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

a. Debt

Preferred stock

Equity

b.Discount rate
Answer #1

1.
=145000*2000*105%/(145000*2000*105%+260000*91+7.2*10^6*65)=0.382460811897106

2.
=260000*91/(145000*2000*105%+260000*91+7.2*10^6*65)=0.0297176446945338

3.
=7.2*10^6*65/(145000*2000*105%+260000*91+7.2*10^6*65)=0.58782154340836

4.
=(145000*2000*105%*(RATE(15*2,5.5%*2000,-2000*105%,2000)*2*(1-21%))+260000*91*(4.4%*100/91)+7.2*10^6*65*(3.1%+1.15*7.5%))/(145000*2000*105%+260000*91+7.2*10^6*65)=10.1591421133763%